# Finite Sample Identification of Partially Observed Bilinear Dynamical Systems

### v1

@@ B.3 Proof of Theorem 2
where we make the simplifying assumption that $T-L+1$ can be divided by $L+1$. (Note that, this goes without loss of generality, and we assume it for the sake of clarity. It can be easily avoided by noting that
$\sum_{t=L}^{T}\tilde{{\bm{u}}}_{t}\tilde{{\bm{u}}}_{t}^{\top}\succeq\sum_{t=L}^{(L+1)\lfloor\frac{T-L+1}{L+1}\rfloor+L-1}\tilde{{\bm{u}}}_{t}\tilde{{\bm{u}}}_{t}^{\top},$
[…]
@@
$\implies\lambda_{\min}(\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}})$ $\geq\inf_{{\bm{v}}_{i}\in\mathcal{N}_{\epsilon}}{\bm{v}}_{i}^{\top}\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}{\bm{v}}_{i}-2\epsilon\|{\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}}\|_{\textup{op}}.$ (B.35)
Hence, in order to lower bound $\lambda_{\min}(\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}})$, we also need an upper bound on $\|{\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}}\|_{\textup{op}}$. This can be done as follows: First, we have
$\operatorname{\mathbb{E}}[\|{\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}}\|_{\textup{op}}]=\operatorname{\mathbb{E}}[\lambda_{\max}(\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}})]$ $\leq\operatorname{\mathbb{E}}[\textup{{tr}}(\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}})]=\textup{{tr}}\big{(}\operatorname{\mathbb{E}}[\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}]\big{)},$
@@
Let :=$\mathcal{E}:=\{\|{\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}}\|_{\textup{op}}\leq\frac{2d_{\tilde{u}}(T-L+1)}{\delta}\}$ denote the event that $\|{\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}}\|_{\textup{op}}$ is upper bounded by a specified threshold. Then, it is straightforward to see that $\operatorname{\mathbb{P}}(\mathcal{E})\geq 1-\delta/2$. This further implies,
$\operatorname{\mathbb{P}}\bigg{(}\lambda_{\min}(\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}})$ $<(T-L+1)(1/2-\eta)\bigg{)}$
($\leq\operatorname{\mathbb{P}}\bigg{(}\big{\{}\lambda_{\min}(\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}})<(T-L+1)(1/2-\eta)\big{\}}\bigcap\mathcal{E}\bigg{)}+\operatorname{\mathbb{P}}(\mathcal{E}^{c}),$
($\leq\operatorname{\mathbb{P}}\bigg{(}\big{\{}\inf_{{\bm{v}}_{i}\in\mathcal{N}_{\epsilon}}{\bm{v}}_{i}^{\top}\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}{\bm{v}}_{i}-2\epsilon\|{\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}}\|_{\textup{op}}<(T-L+1)(1/2-\eta)\big{\}}\bigcap\mathcal{E}\bigg{)}+\delta/2,$
@@
$\operatorname{\mathbb{P}}\bigg{(}\lambda_{\min}(\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}})$ $<(T-L+1)/2$
@@
we repeat the Step 3) in the proof of Theorem 2 with :=$\mathcal{E}:=\{\|{\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}}\|_{\textup{op}}\leq(T-L+1)d_{\tilde{u}}\}$ and using $\epsilon=\frac{1}{4d_{\tilde{u}}}$-net covering argument to get the following persistence of excitation result: For all $\delta\in(0,1)$, the event,
@@
$T-L+1\geq 32(L+1)\Xi\big{(}\log\big{(}{(L+1)}/{\delta}\big{)}+d_{\tilde{u}}\log\big{(}1+8d_{\tilde{u}}\big{)}\big{)}.$ (B.45)
@@ C.3 Proof of Proposition 2
=$\qquad\overset{(a)}{=}\sum_{t=1}^{T-L}(\theta^{\top}\tilde{{\bm{u}}}_{t+L})(\lambda^{\top}{\bm{C}})\left(\prod_{\ell=1}^{L-1}({\bm{u}}_{t+\ell}\circ{\bm{A}})\right)\sum_{s=0}^{t-1}\left(\prod_{k=s+1}^{t-1}({\bm{u}}_{k}\circ{\bm{A}})\right)({\bm{B}}{\bm{u}}_{s}+{\bm{w}}_{s})$
@@
=$\qquad\overset{(b)}{=}\sum_{t=0}^{T-L-1}\sum_{s=0}^{t}\underbrace{(\theta^{\top}\tilde{{\bm{u}}}_{t+1+L})(\lambda^{\top}{\bm{C}})\left(\prod_{\ell=1}^{L-1}({\bm{u}}_{t+1+\ell}\circ{\bm{A}})\right)}_{:={\bm{M}}({\bm{u}}_{t+2},\dots,{\bm{u}}_{t+1+L})}\underbrace{\left(\prod_{k=s+1}^{t}({\bm{u}}_{k}\circ{\bm{A}})\right)}_{:={\bm{N}}({\bm{u}}_{s+1},\dots,{\bm{u}}_{t})}({\bm{B}}{\bm{u}}_{s}+{\bm{w}}_{s})$

### v2

@@ B.3 Proof of Theorem 2
where we make the simplifying assumption that TL+1T-L+1 can be divided by $L+1$. (Note that, this goes without loss of generality, and we assume it for the sake of clarity. It can be easily avoided by noting that _t=L^T ~u_t ~u_t^⊤⪰∑_t=L^ (L+1)T-$L+1$ + L-1 ~u_t ~u_t^, where $\lfloor\cdot\rfloor$ denotes the floor operator. We can then analyze everything with $T_{0}=(L+1)\lfloor\frac{T-L+1}{L+1}\rfloor+L-1$, and note that $T-L\leq T_{0}\leq T$.)
@@
[…]
@@
Let $\mathcal{E}:=\{\|{\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}}\|_{\textup{op}}\leq\frac{2d_{\tilde{u}}(T-L+1)}{\delta}\}$ denote the event that $\|{\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}}\|_{\textup{op}}$ is upper bounded by a specified threshold. Then, it is straightforward to see that $\operatorname{\mathbb{P}}(\mathcal{E})\geq 1-\delta/2$. This further implies,
[…]
$\leq\operatorname{\mathbb{P}}\bigg(\big\{\lambda_{\min}(\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}})<(T-L+1)(1/2-\eta)\big\}\bigcap\mathcal{E}\bigg)+\operatorname{\mathbb{P}}(\mathcal{E}^{c}),$
$\leq\operatorname{\mathbb{P}}\bigg(\big\{\inf_{{\bm{v}}_{i}\in\mathcal{N}_{\epsilon}}{\bm{v}}_{i}^{\top}\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}{\bm{v}}_{i}-2\epsilon\|{\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}}\|_{\textup{op}}<(T-L+1)(1/2-\eta)\big\}\bigcap\mathcal{E}\bigg)+\delta/2,$
@@
[…]
@@
we repeat the Step 3) in the proof of Theorem 2 with $\mathcal{E}:=\{\|{\tilde{{\bm{U}}}^{\top}\tilde{{\bm{U}}}}\|_{\textup{op}}\leq(T-L+1)d_{\tilde{u}}\}$ and using $\epsilon=\frac{1}{4d_{\tilde{u}}}-$net covering argument to get the following persistence of excitation result: For all $\delta\in(0,1)$, the event,
@@
$).\displaystyle T-L+1\geq 32(L+1)\Xi\big(\log\big({(L+1)}/{\delta}\big)+d_{\tilde{u}}\log\big(1+8d_{\tilde{u}}\big$$\big)$. (B.45)
@@ C.3 Proof of Proposition 2
$\qquad\overset{(a)}{=}\sum_{t=1}^{T-L}(\theta^{\top}\tilde{{\bm{u}}}_{t+L})(\lambda^{\top}{\bm{C}})\left(\prod_{\ell=1}^{L-1}({\bm{u}}_{t+\ell}\circ{\bm{A}})\right)\sum_{s=0}^{t-1}\left(\prod_{k=s+1}^{t-1}({\bm{u}}_{k}\circ{\bm{A}})\right)({\bm{B}}{\bm{u}}_{s}+{\bm{w}}_{s})$
@@
$\qquad\overset{(b)}{=}\sum_{t=0}^{T-L-1}\sum_{s=0}^{t}\underbrace{(\theta^{\top}\tilde{{\bm{u}}}_{t+1+L})(\lambda^{\top}{\bm{C}})\left(\prod_{\ell=1}^{L-1}({\bm{u}}_{t+1+\ell}\circ{\bm{A}})\right)}_{:={\bm{M}}({\bm{u}}_{t+2},\dots,{\bm{u}}_{t+1+L})}\underbrace{\left(\prod_{k=s+1}^{t}({\bm{u}}_{k}\circ{\bm{A}})\right)}_{:={\bm{N}}({\bm{u}}_{s+1},\dots,{\bm{u}}_{t})}({\bm{B}}{\bm{u}}_{s}+{\bm{w}}_{s})$